derivative y = arctan4(3x5)


1
Expert's answer
2022-04-14T13:58:33-0400

Remind that the derivative of the composition of functions"f(g(x))" is: "f'_x(g(x))=f'_g(g(x))g'(x)". Thus, we get: "y'=\frac{4(\arctan^3(3x^5))(3x^5)'}{1+9x^{10}}=\frac{60x^4\arctan^3(3x^5)}{1+9x^{10}}."

The answer is: "y'=\frac{60x^4\arctan^3(3x^5)}{1+9x^{10}}"